OpenStudy (anonymous):

What decimal is equivalent to 7 3/250? 7 .12 B) 7. 012 C) 7.325 D) 7.3025

OpenStudy (unklerhaukus):

\[7\frac3{250}=7+\frac6{500}=7+\frac{12}{1000}=7+\frac{1.2}{100}\]

OpenStudy (anonymous):

thank you some much for shoeing me the steps :)

OpenStudy (unklerhaukus):

\[=7+\frac{0.12}{10}=7+\frac{0.012}{1}=7+{0.012}\]